A researcher wishes to estimate the proportion of adults who have​ high-speed Internet access. What size sample should be obtained if she wishes the estimate to be within 0.05 with 90​% confidence if

​(a) she uses a previous estimate of 0.42​?

​(b) she does not use any prior​ estimates?

a)n=

​(Round up to the nearest​ integer.)

​(b)n=

​(Round up to the nearest​ integer.)
Answer #1

Ans:

Margin of error,E=0.05

confidence level=0.90,so z=1.645

a)prior estimate=0.42

Sample size required,n=1.645^2*0.42*(1-0.42)/0.05^2=264

b)No prior estimate,so consider p=0.5

Sample size required,n=1.645^2*0.5*(1-0.5)/0.05^2=271